Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 113     Entry time: 30 Mar 2023, 02:37
Author: Jörg Hallmann 
Group: MID 
Subject: shift summary March 29th 

achievments: alignment of the SDL (both branches)

measuring of samples & calibrants

 

issues:

- broken laser interlock - cannot be set - PRC is informed and contacted SRP - repair expected in the morning

- ScannerX broken - morning shift might contact DOC or expert

- EPIXs do not send data after rearrangement of the modules - expert consulation might be necessary

ELOG V3.1.4-7c3fd00